### v4

# SWAA: Sliding Window Attention Adaptation for Efficient Long-Context LLMs Without Pretraining

Abstract
The quadratic complexity of self-attention in Transformer-based Large Language Models (LLMs) renders long-context inference prohibitively expensive. While Sliding Window Attention (SWA), the simplest sparse attention pattern, offers a linear-complexity alternative, naively applying it to models pretrained with Full Attention (FA) causes catastrophic long-context performance collapse due to the training-inference mismatch. To address this, we propose Sliding Window Attention Adaptation (SWAA), a plug-and-play toolkit of recipes that adapt FA models to SWA without costly pretraining. SWAA systematically combines five strategies: (1) applying SWA only during prefilling; (2) preserving “sink” tokens; (3) interleaving FA/SWA layers; (4) chain-of-thought (CoT); and (5) fine-tuning. Our experiments demonstrate that while individual methods are insufficient, specific synergistic combinations can effectively recover original long-context capabilities. After further analyzing performance-efficiency trade-offs, we identify recommended SWAA configurations for diverse scenarios, which achieve 30% to 100% speedups for long-context LLM inference with acceptable quality loss. Our code is available at github.
SWAA: Sliding Window Attention Adaptation for Efficient Long-Context LLMs Without Pretraining
1 Introduction
Transformer-based Large Language Models (LLMs) Vaswani et al. (2017) demonstrate remarkable capabilities, but their self-attention scales quadratically with the input sequence length, making long context processing inefficient. Sliding Window Attention (SWA), the most straightforward and widely adopted sparse attention pattern, which restricts each token’s attention to a fixed-size local window, reduces the computational complexity to linearity, along with some other benefits (see Appendix A).
Despite its simplicity and ease of use, significant challenges remain in leveraging SWA to accelerate LLM inference. On one hand, due the performance degradation brought by training-inference attention pattern mismatch, to apply SWA to LLMs, most typical solutions involve training a model with SWA from scratch, but they are prohibitively costly and cannot match the performance of state-of-the-art full-causal-attention models like Qwen3 Team (2025b), mainly due to the inability to reproduce pretraining data. On the other hand, training-free methods like streaming attention Xiao et al. (2024) can stabilize LLM outputs by retaining “sink tokens” while applying SWA, which greatly improve efficiency, but inevitably suffer from severe long-context performance degradation possibly due to the inaccessibility of distant tokens’ information Xiao (2025). This motivates a critical, unexplored question: Can we adapt a full-attention model to sliding window attention at low cost while maintaining long-context performance?
We answer Yes to this question by proposing Sliding Window Attention Adaptation(SWAA), a set of recipes for adapting FA-pretrained models to SWA, which requires neither costly pretraining nor modifications to the standard Transformer architecture. Specifically, it systematically combines five practical and composable methods:
  1. 1.
    Full Attention (FA) Decode: applying SWA only during the prefilling stage while switching back to full attention for decoding.
  2. 2.
    Keep First $k$ Tokens: explicitly preserving attention to the first $k$ “sink” tokens.
  3. 3.
    Interleaving FA/SWA layers: mixing full-attention and SWA layers (e.g., applying SWA to half of the layers).
  4. 4.
    Chain-of-Thought (CoT): enforcing an explicit "thinking" process during decoding.
  5. 5.
    Fine-tuning with SWA: lightweight SWA-aware supervised fine-tuning on long-context data.
Among these, FA Decode is a novel method we introduce. Keep First $k$ Tokens and FA/SWA Interleaving have been proven effective in prior work Xiao et al. (2024); Team (2024a); Zhang et al. (2024), while CoT and fine-tuning are common LLM techniques. However, the optimal strategy for combining these methods to achieve effective SWA adaptation remains unexplored.
Therefore, in our experiments, we evaluate SWAA on Qwen3 Team (2025b) and Llama3.1 Team (2024b) across several long-context benchmarks, measuring both performance and efficiency under a wide range of SWAA configurations. First, we find that each method makes a distinct contribution, but no single ingredient suffices to make SWA competitive with full attention. Second, we show that specific synergistic combinations of methods can recover a large fraction of the original long-context performance. Third, we analyze the performance-efficiency trade-offs of different SWAA recipes and identify some recommended configurations suitable for different deployment scenarios, which can achieving nearly 100% speedup with 90% accuracy retention in efficiency-first scenarios, or 30% speedup with nearly 100% accuracy retention in quality-first scenarios. This shows SWAA is a flexible toolkit beyond a fixed strategy.
Our key contributions are:
  1. 1.
    We perform the first systematic study on adapting FA-pretrained LLMs to SWA without pretraining, revealing novel insights about how SWA impacts LLMs and providing a foundation for future research in efficient sparse attention.
  2. 2.
    We propose SWAA, a set of practical SWA adaptation recipes that offer a robust performance-efficiency balance for various use cases, accelerating LLM inference from the bottom level.
  3. 3.
    We implement our methods with Flash-Attention Dao (2024) and vLLM Kwon et al. (2023), making it plug-and-play and user-friendly for practical deployment.
2 Related Works
The $O(N^{2})$ complexity of self-attention in Transformers Vaswani et al. (2017) has spurred a wide field of research about more efficient language model architectures. Among the two most popular technological routes are sparse attention and linear attention.
2.1 Sparse Attention
Our work falls in this category. Sliding Window Attention (SWA) represents the most basic form of local sparse attention, yet its performance is inherently limited. Therefore, model architectures such as Longformer Beltagy et al. (2020), BigBird Zaheer et al. (2020), and RATTENTION Wang et al. (2025) combine local SWA on most tokens with special global attention on specific tokens to create a more powerful, albeit still sparse, pattern. Popular LLMs like Gemma2 Team (2024a) adopt SWA in half of their layers to balance the efficiency of SWA and performance of FA. Sliding Window Attention Training (SWAT) Fu et al. (2025b) introduces architectural changes, such as sigmoid activation and balanced position embeddings, to stabilize SWA performance. More advanced methods like Deepseek-sparse-attention Yuan et al. (2025); DeepSeek-AI (2025b), although achieving excellent quality, involve more complicated implementation and optimization due to semantic-aware attention operations (e.g., selecting the most important tokens based on attention weights). Nevertheless, nearly all aforementioned methods require pretraining with a specific sparse pattern, which is costly and fails to leverage the advantages of existing pretrained models.
LightTransfer Zhang et al. (2024) is a promising attempt at adapting existing models to SWA without pretraining, which has the same motivation as ours. But it may generalize poorly across model families (see Appendix F).
2.2 Linear Attention
An alternative approach involves reformulating the attention mechanism entirely to achieve linear, $O(N)$, complexity. This includes methods such as RNN-like linear attention transformers Katharopoulos et al. (2020); Peng et al. (2023); Sun et al. (2023) and structured state-space models (SSMs) like Mamba Gu and Dao (2023). Many works such as Jamba and Nemotron-FlashLieber et al. (2024); Linsong Chu and others (2024); Team et al. (2025); Fu et al. (2025a) interleave linear attention layers with traditional attention layers to create hybrid model structures. While promising, they represent a fundamental architectural departure from the standard Transformer, and their performance is generally weaker than traditional Transformer-based LLMs.
3 Candidate Methods for SWA Adaptation
As previously established, naive application of SWA leads to severe long-context performance degradation. Therefore, we investigate five methods that can potentially facilitate SWA adaptation from distinct perspectives. However, every method except fine-tuning has some drawbacks to LLM inference efficiency, as discussed in Appendix A. Therefore, although these methods are not mutually exclusive, indiscriminately adopting all of them may not achieve the optimal performance-efficiency trade-off. Instead, we must evaluate various combinations to identify the optimal trade-off.
3.1 Full Attention Decode
This method applies SWA only to the prefilling stage. During the decoding (auto-regressive generation) stage, each token still employs full attention, allowing access to all previous tokens in the context. The resulting attention mask is depicted in Figure 1(a).
This novel approach, first proposed in our study, is inspired by human reading comprehension: humans typically scan a passage casually (prefilling) before thinking deeply to formulate an answer (decoding) for a specific problem. We term this strategy "reading casually, thinking carefully." In our design, the SWA-constrained prefilling stage corresponds to casual reading, while the full-attention decoding stage enables careful thinking. This analogy also suggests that Chain-of-Thought (CoT) during decoding may be particularly beneficial, as extended generation (i.e. "thinking") could compensate for the insufficient contextual information gathered during the prefilling stage.
Figure 1: (a) Attention mask for FA Decode. SWA is used for prompt tokens (prefill), and full attention is used for generated tokens (decode). (b) Attention mask for SWA combined with Keep First $k$ Tokens.
3.1.1 Keep First k Tokens
Streaming attention Xiao et al. (2024) demonstrates that models pretrained with full attention allocate a disproportionate amount of attention to the initial tokens ("attention sink"), so it permanently retains the attention to these "sink" tokens while using SWA, which successfully maintains the stability of the attention distribution and the model’s output. Our method is basically the same: as shown in Figure 1(b), any subsequent token can attend to its local window and the initial $k$ tokens.
Crucially, we extend this approach further. Streaming attention operates only at the KV cache level; specifically, the KV cache of the sink tokens is retained during decoding, while the prefilling stage is not modified or accelerated at all. In contrast, we directly customize the Flash-Attention-2 Dao (2024) kernel to implement such attention mask, thereby accelerating prefilling via SWA as well, and eliminating the need to modify KV cache.
3.2 Interleaving Layers
This method retains full attention on a subset of layers while applying SWA to the remainder, providing a simple hybrid mechanism to balance the performance of full attention with the efficiency of pure SWA. A common strategy involves designating one in every $n$ layers to use full attention (e.g., layers 0, 2, 4, …retain full attention, while all others use SWA). For example, Gemma-2 Team (2024a) uses SWA only for layers [1, 3, 5, …], and Gemma-3 Team (2025a) uses SWA only for layers [5, 11, 17, …].
However, for an FA-pretrained model, layers may exhibit distinct behaviors, suggesting it may not be optimal to simply assign SWA to layers [1, 3, 5, …]. Some methods like LightTransfer Zhang et al. (2024) have explored more fine-grained layer selection, but we find it is not consistently superior in practice (see Appendix F). Therefore, for simplicity, we still choose the simplest fixed-interval layer selections in our experiments, such as [0, 2, 4, …] and [1, 3, 5, …].
3.3 Chain-of-Thought
Chain-of-Thought (CoT) Wei et al. (2022) is a widely used technique for improving model accuracy via reasoning. With the advent of "thinking" models, such as DeepSeek-R1 DeepSeek-AI (2025a), CoT has evolved from a prompting strategy to an intrinsic LLM capability. However, whether CoT has a specific impact in SWA scenarios remains uninvestigated. We explore this by comparing a thinking model with its non-thinking variant, e.g., Qwen3-4B-Thinking and Qwen3-4B-Instruct, to verify the effect of CoT on SWA adaptation, which can produce more notable differences compared to simply adding CoT prompting to the same model.
3.4 Fine-tuning
Fine-tuning the model with SWA enabled is the most natural way to mitigate training-inference mismatch. Meanwhile, the training data must be long-context, so that SWA actually works.
However, most available long-context datasets only contain brief ground-truth answers, lacking the reasoning steps required for fine-tuning a "thinking" model. Since our goal is to restore the model’s original capabilities under SWA rather than teach it new ones, we adopt an approach similar to self-distillation Yang et al. (2024). Specifically, we utilize the original full-attention model to generate new answers for the dataset’s questions, and these generated answers are then filtered for correctness using GPT-5-Mini OpenAI (2025), to construct our training dataset. For each question, we sample 4 answers with temperature 1 to enhance diversity, because we find it is slightly better than generating only one answer.
Table 1: The accuracy (%) of Qwen3-4B on LongMemEval, LongBench-V2, and Ruler (without SFT)
No. Window FA layers Keep FA LongMem LB-V2 Ruler
size first decode Think Non-T Think Non-T Think Non-T
0 Full [] 0 False 73.0 62.0 34.6 35.2 85.6 92.8
1 2k [] 0 False 3.2 11.0 9.4 25.8 0.0 0.0
2 8k [] 0 False 13.2 19.8 15.1 22.1 0.0 0.0
3 2k [] 10 False 16.0 15.6 7.7 25.8 0.0 0.0
4 2k [] 0 True 11.8 14.2 26.2 25.2 0.2 0.0
5 2k [1, 3, 5, …] 0 False 13.4 18.4 12.1 23.5 0.0 0.0
6 8k [] 0 True 26.2 25.0 22.8 25.5 0.0 0.0
7 2k [] 10 True 38.2 20.6 25.8 25.2 1.0 0.0
8 2k [] 100 True 50.0 17.8 24.2 26.5 3.2 0.4
9 2k [] 1000 True 50.0 20.2 23.8 25.2 3.2 0.2
10 2k [0, 2, 4, …] 10 False 17.0 14.8 19.8 29.2 0.0 0.0
11 2k [0, 2, 4, …] 0 True 32.2 26.0 23.8 29.9 1.0 2.0
12 2k [1, 3, 5, …] 10 False 25.8 36.4 21.1 29.9 0.0 0.0
13 2k [1, 3, 5, …] 0 True 59.2 34.8 28.5 26.8 22.8 32.8
14 4k [] 10 True 38.0 24.4 27.9 27.5 1.6 2.2
15 8k [] 10 True 49.2 35.2 36.7 30.2 3.2 6.2
16 2k [0, 2, 4, …] 10 True 36.0 17.2 30.5 30.2 10.0 3.0
17 2k [1, 3, 5, …] 10 True 65.0 53.6 35.4 28.6 59.2 75.8
18 2k [1, 3, 5, …] 100 True 68.8 50.6 35.6 31.8 58.0 78.8
19 2k [1, 5, 9, …] 10 True 53.2 31.4 29.6 26.4 9.8 26.4
20 2k [1, 9, 17, …] 10 True 36.4 18.8 28.6 25.4 0.6 0.2
21 2k [3, 7, 11, …] 10 True 54.2 34.6 27.7 26.0 1.6 4.2
22 4k [1, 3, 5, …] 100 True 73.0 54.2 33.4 31.2 65.4 83.6
23 8k [1, 3, 5, …] 100 True 71.6 56.6 35.4 31.5 67.2 86.0
Table 2: The accuracy (%) of Qwen3-4B on LongMemEval, LongBench-V2, and Ruler (with SFT)
No. Window FA layers Keep FA LongMem LB-V2 Ruler
size first decode Think Non-T Think Non-T Think Non-T
0 Full [] 0 False 74.6 63.4 37.9 34.9 88.2 90.6
1 2k [] 0 False 18.8 23.8 7.4 30.9 0.0 0.0
2 2k [] 100 False 15.6 $\backslash$ 6.0 $\backslash$ 0.0 $\backslash$
3 2k [] 0 True 57.9 42.0 29.2 30.2 0.4 0.8
4 2k [1, 3, 5, …] 0 False 63.6 54.6 29.5 31.9 59.0 62.2
5 4k [] 0 True 62.6 $\backslash$ 32.9 $\backslash$ 0.2 $\backslash$
6 2k [] 10 True 56.7 $\backslash$ 28.9 29.2 9.2 14.4
7 2k [] 100 True 62.2 42.6 29.2 30.5 13.4 20.4
8 2k [1, 3, 5, …] 0 True 73.2 58.8 38.3 34.6 74.0 85.8
9 2k [0, 2, 4, …] 0 True 66.0 $\backslash$ 31.5 $\backslash$ 8.0 $\backslash$
10 2k [1, 5, 9, …] 0 True 68.8 47.0 32.0 32.2 7.6 28.0
11 2k [1, 3, 5, …] 100 True 73.2 61.4 37.2 33.9 72.8 89.2
4 Experiment Setup
We organize our experiments around three research questions:
RQ1:
Is SWA adaptation feasible without any additional training?
RQ2:
How much does fine-tuning with SWA improve performance?
RQ3:
Which SWAA configurations achieves the optimal performance-efficiency trade-offs?
4.1 Models
Our primary experiments use Qwen3-4B-Thinking and Qwen3-4B-Instruct Team (2025b). The Thinking variant enforces chain-of-thought (CoT) style reasoning, whereas the Instruct variant usually just answers briefly. To ensure generality, we additionally evaluate Qwen3-30B-A3B-Thinking, Qwen3-30B-A3B-Instruct Team (2025b), and Llama3.1-8B-Instruct Touvron et al. (2023), as shown in Appendix C.
All models are served with vLLM in float16 precision using a batch size of 64, with our customized Flash-attention-2 kernel to support Keep First and FA Decoding. We use greedy decoding (temperature =0=0) for all evaluations. In preliminary experiments, we observed that vLLM yields slightly lower (about 1% to 5%) scores than HuggingFace Transformers due to precision-related discrepancies.
4.2 Evaluation Dataset
SWA is identical to full attention when the context length is within the window size. Even if the model is fine-tuned, we can simply disable the LoRA adapters for short prompts to get the same response as the original model. Therefore, our evaluations focus on long-context benchmarks, as re-evaluating models on standard short-context benchmarks (e.g., MMLU Hendrycks et al. (2021), GPQA Rein et al. (2023)) is completely unnecessary.
We select LongMemEval Wu et al. (2024), a benchmark consisting of various types of long-context QA tasks with moderate difficulty, as our main benchmark, although it is originally designed for agent memory system evaluation. Its context length is controllable by selecting a specific number of chat sessions to concatenate as the context from a pool of hundreds of sessions (a session contains the chat history between user and assistant within a day). To create a moderately difficult and discriminative evaluation, we construct LongMemEval_24k by sampling 10 sessions, resulting in 500 samples ranging from 16k to 32k with an average context length of 24k.
We also experiment on LongBench-V2 Bai et al. (2024b) and Ruler Hsieh et al. (2024), which are most commonly used long-context benchmarks, for additional validation of generalizability. For LongBench V2, we retain only the samples whose context length is under 128k due to GPU memory limitations; thus, 311 of 500 samples are kept. For Ruler, we choose the Multi-Query task, which contains 500 samples, and control the context length to 128k (counted by Qwen3 tokenizer).
To judge the correctness for accuracy, we use LLM-as-judge with GPT-5-Mini OpenAI (2025) for LongMemEval and exact-match for LongBench-V2 and Ruler.
However, we find other LongBench-V2 is too difficult for 4B-level models, and Ruler’s task types are not diverse enough (see details in Appendix B), so our analysis will more focus on the clearer results of LongMemEval.
4.3 Training Details
For the fine-tuning dataset, we initially considered LongAlign Bai et al. (2024a), a widely used long-context fine-tuning dataset for adapt a regular-length model to long-context tasks. However, since its sample count ($\sim$10,000) is insufficient, we incorporate an additional 6,000 samples from Fusang-v1-long Pan (2024), a more comprehensive corpus of over 40,000 long-context samples that includes LongAlign as a subset.
We perform SWA-aware fine-tuning using LoRA Hu et al. (2022). Unless otherwise noted, we use rank $r=16$ and $\alpha=128$, and apply LoRA only to the query, key, and value projection modules. We adopt this parameter-efficient setting because full-parameter fine-tuning often leads to overfitting and degradation of the model’s original capabilities in our preliminary experiments. We use a learning rate of 1e-4 with a cosine decay schedule. Models are fine-tuned for a single epoch on the sampled long-context dataset since we observe no meaningful gains from additional epochs (see Appendix E). Training of each SWAA configuration takes approximately 12 hours on an 8*H20 GPU server for Qwen3-4B and 30 hours for Qwen3-30B-A3B.
5 Experiment Results
5.1 SWA Adaptation Without Fine-tuning
We first investigate training-free SWA adaptation. Table 1 reports LongMemEval_24k accuracy for Qwen3-4B-Thinking ("think") and Qwen3-4B-Instruct ("non-think") under different combinations of SWAA components. In most settings, we use an aggressive 2k window to amplify the impact of SWA. The configurations are grouped by the number of methods applied (0, 1, 2, or 3 of Interleaving Layers, Keep First and FA Decode). Rows 0 (original model) and 1 (naive SWA) serve as upper and lower baselines, respectively. In the column "FA layers", the value records which layers use full attention, and [] means no layers. In the column "Keep first", the value is $k$ in Keep First $k$ Tokens. From the results of LongMemEval, we find that:
Naive SWA is not viable. For LongMemeval, naively replacing FA with a 2k sliding window attention (row 1) drops accuracy significantly to $3.2$ and $11.0$, respectively. Even with an 8k window (row 2), accuracy only recovers to $13.2$ and $19.8$, far below the FA baseline.
Individual methods provide gains but fail to close the performance gap. Each method—Keep First, FA Decode, or Interleaving Layers—improves over naive SWA (rows 3–6), yet each alone recovers only a small fraction of the FA gap and remains well below the baseline. In short, no single method is sufficient.
Combinations exhibit strong synergy. Recipes that combine multiple methods deliver large gains: (1) FA Decode + Keep First $k$ substantially improves over naive SWA (rows 7–9), recovering approximately half to two-thirds of the performance drop for the thinking model as $k$ increases. However, increasing $k$ from 100 to 1000 yields little improvement, indicating that $k$ does not need to be exceedingly large. (2) Interleaving Layers + FA Decode is markedly stronger (row 13), recovering most of the gap for the thinking model. (3) FA Decode + Interleaving Layers + Keep First $k$ perform best (rows 18). The thinking model recovers close to 90% of the FA gap even at 2k window.
CoT synergizes with FA Decode. Under recipes that include FA Decode, the thinking model consistently benefits more than the non-thinking model (rows 13 and 18), suggesting that CoT synergizes with FA Decode: preserving global attention at decoding time enables longer reasoning traces to capitalize on context processed by SWA, confirming our hypothesis in Section 3.1.
Sliding window size affects, but is not the decisive role. With FA Decode + Keep First $k$, accuracy improves as the window increases (rows 7, 14, 15), though benefits are modest until 8k. When added with interleaving FA layers, moving from a 2k to a 4k window is enough to close the remaining gap on the thinking model (row 22 matches the FA baseline), indicating that FA Decode and Interleaving are the dominant drivers, and larger windows mainly smooth residual error.
SWA/FA layer selection has large impacts. Selecting less layers for FA, e.g. only one fourth or eighth (row 19, 20), though more efficient, will significantly decrease the improvement brought by Interleaving Layers. More importantly, for Qwen3-4B, configuring odd-numbered layers with full attention is significantly better than even-numbered layers (row 11, 13, 16, 17). However, surprisingly, this result is reversed for Qwen3-30B-A3B and Llama3.1-8B-Instruct (row 10 and 12 in Table 4 and 5). This suggests that layer functionalities differ across model families and sizes, necessitating model-specific layer selection strategies, as discussed in Section 3.2.
For LongBench V2, due to the high difficulty, the performance is generally poor. Some scores are even below the random guessing baseline (25%). However, despite narrower performance gaps between configurations, the overall accuracy trend still aligns with that of LongMemEval, except for few outliers, which may be attributed to the inherent randomness of multiple-choice questions.
For Ruler, as it involves only the needle-in-a-haystack task, performance varies drastically between configurations. The accuracy of the thinking model is even always lower than the non-thinking one. And the accuracy is hardly improved until we let at least half layers use FA and enable FA decoding at the same time.
Therefore, we answer RQ1: adapting an FA LLM to SWA is feasible even without any training. But it requires specific combinations of at least 2 methods, which could be less efficient for inference.
5.2 SWA Adaptation With Fine-tuning
We next evaluate SWA-aware supervised fine-tuning, which is expected to provide higher improvement. Table 2 reports LongMemEval_24k accuracy after SFT under various SWAA configurations. The original full-attention model is also fine-tuned as a baseline (row 0). Since training is relatively time-consuming, we only test a representative subset of configurations. From the results of LongMemEval, our findings are as follows:
Fine-tuning substantially lifts all SWA configurations. Comparing all the fine-tuning results with non-training ones, it is clear that fine-tuning consistently provides great improvement. However, simply fine-tuning with naive SWA remains insufficient, only achieving 18.8% and 23.8% accuracy (row 1).
FA Decode and Interleaving Layers emerge as dominant components. After SFT, FA Decode and Interleaving Layers provide the largest gains. Using FA Decode alone (row 3) or using Interleaving Layers alone (row 4) both get high accuracy. And combining both (row 8) further boosts performance to $73.2$ (think) and $58.8$ (non-think), nearly matching the full-attention SFT baseline in row 0.
Keep First becomes optional rather than essential. Before fine-tuning, Keep First is crucial for stability under SWA. But after SFT, it only provides minor additional improvement. With FA Decode, adding $k=100$ (row 7) improves over $k=0$ (row 3) only 4.5%, and if further combining FA layers, it almost offers no improvements (row 11 and row 8).
Effect of sliding window size. Row 3 and 5 shows that increasing the window from 2k to 4k with FA Decode improves thinking-model accuracy from $57.9$ to $62.6$. This mirrors the non-SFT trend that larger windows help, but the dominant improvements still come from FA Decode and Interleaving Layers.
The improvement on LongBench-V2 is mostly consistent with that on LongMemEval. While for Ruler, SFT also brings improvements, but configuring odd-numbered layers with full attention has a decisive effect that is far greater than other methods. Notably, long reasoning has no effect or even a negative effect on Ruler, which may also be the reason for the little improvement of FA decoding.
Thus, regarding RQ2, we conclude that fine-tuning yields substantial performance restoration, provided we apply FA Decode, Interleaving Layers, or a combination thereof, while Keep First becomes optional. And the improvement brought by SFT under each configuration varies significantly, implying that a configuration optimal for the training-free setting is not necessarily optimal after SFT, and vice versa.
Figure 2: Accuracy and inference time of each configuration of Qwen3-4B on LongMemEval
5.3 Performance–efficiency Trade-offs and Recommended Recipes
Although integrating more methods can typically achieve higher accuracy, it introduces more overhead, indicating that the efficiency of each recipe must also be evaluated. To assess the performance-efficiency trade-off of different SWAA configurations, we evaluate time-to-first-token (TTFT), time-per-output-token (TPOT), total throughput, and average running time per request. Concretely, we benchmark Qwen3-4B-Thinking on a single H100 GPU using vLLM’s bench_serve utility Kwon et al. (2023) with random input data and 100 total requests. The prompt length and output length are set to 128k and 512 tokens, respectively, representing a typical long-context QA setting.
To visualize the performance-efficiency trade-off, Figure 2 plots each configuration’s accuracy on LongMemEval_24k Wu et al. (2024) against its average running time, while detailed TTFT, TPOT, and throughput statistics for each configuration are provided in Appendix D. We draw a line between the full-attention point and the naive-SWA point as a baseline curve: configurations above this line offer a better accuracy-latency balance intuitively. For configurations with nearly identical time costs, we display only the one with the highest accuracy. Since Keep First $k$ has negligible impact on runtime (Appendix D), all plotted configurations fix $k=10$.
Table 3: Recommended SWA adaptation recipes for different needs and scenarios: whether you want to train, use a thinking model, or prefer efficiency or quality. means optional.
Train Thinking Priority FA Decode Inter-leaving Keep First
No No Any
No Yes Efficiency
No Yes Quality
Yes Any Efficiency
Yes Any Quality
From Figure 2, we observe that many configurations in Figure 2 achieve a clearly better performance-efficiency ratio than baselines. For example, after SFT, (1) enable either FA decoding or Interleaving Layers can enhance the inference speed by almost 100%, while maintaining nearly 90% of the original accuracy; (2) while combining them can achieve nearly 100% original accuracy maintenance and over 30% acceleration; (3) By controlling the number of layers using FA, efficiency and quality can be achieved between that of (1) and (2). And for the thinking model, more points lie above the baseline curve compared to non-thinking, indicating that CoT generally has a positive effect on improving the performance-efficiency ratio of SWAA.
Thus, we finally answer RQ3: many SWAA configurations all reach excellent performance-efficiency trade-off, but there is no single metric to quantify such trade-off to decide the globally optimal one. We therefore summarize recommended SWA adaptation recipes tailored to various deployment scenarios in Table 3. And we must note that specific parameters should be flexibly set to meet application-specific requirements, without the need to follow our experimental parameters (e.g., a 2k window, $k=10$). For example, you can increase the window size to 4k or $k$ to 128 for higher accuracy and acceptable additional overhead.
6 Conclusion
In this work, we validate the feasibility of adapting full-attention pretrained LLMs to Sliding Window Attention (SWA) for better efficiency, offering a cost-effective alternative, SWAA, that avoids training sparse-attention models from scratch. By systematically deconstructing the adaptation process, we identify that the catastrophic degradation observed in naive implementations can be effectively mitigated through synergistic combinations of auxiliary methods. Our extensive experiments across Qwen and Llama families demonstrate that while trade-offs between computational overhead and model performance are inevitable, optimized configurations can achieve excellent performance-efficiency balance to effectively accelerate LLM inference.
7 Limitations
We hypothesize that the ideal reasoning trajectory of the model adapted to SWA should be longer than the original model, to compensate for the information loss caused by SWA. That means, using the answers generated by the original model as fine-tuning data may not be the optimal training method. Rather, RL methods like GRPO Shao et al. (2024) might further help the model adapted to SWA by learning a more adaptive reasoning trajectory. However, we did not experiment with them since they are too time-consuming and unstable.
We have not yet implemented the KV cache eviction (or overwriting) mechanism when using SWA; that is, although the speed is improved, the memory usage is actually unchanged. That is because not all of our SWAA configurations support discarding early KV cache.
Appendix A SWA’s Benefits and Each Method’s Drawbacks
SWA reduces the computational complexity to $O(N\cdot W)$, where $W$ is the window size. The benefits are threefold: (1) SWA reduces the computational load, (2) conserves GPU memory by limiting the required Key-Value (KV) cache, and (3) enhances KV cache reusability beyond traditional prefix caching, since a token’s state is independent of tokens outside its local window. However, there is no free lunch—each method of SWAA has some drawbacks, impairing the benefits brought by SWA to varying degrees.
FA Decode presents two primary drawbacks: (1) the benefits apply only to prefilling, while decoding speed is not accelerated as it utilizes full attention, and (2) the GPU memory required for the KV cache is not reduced, as the KV cache for the full context must be retained for decoding. In practice, however, many distributed LLM services have to recompute the KV cache of the entire chat history because storing and loading the KV cache complicates engineering systems, making prefilling occurs more frequently than expected, thereby amplifying the advantage of this method.
Keep First introduces very minor computational overhead, but it complicates efficient KV cache reuse. Due to positional encoding, a token’s KV state depends on its position relative to the initial $k$ tokens, hindering simple cache reuse across different requests. A position encoding separation or offsetting mechanism may be needed.
Interleaving Layers introduces the most significant overhead, as only a subset of layers benefits from the computational savings of SWA. Furthermore, the GPU memory required for the KV cache is not reduced for the full-attention layers. Additionally, this method negates the KV cache reusability advantage of SWA, as the existence of full-attention layers violates the independence of the KV cache beyond the local window.
CoT will greatly increase the generation length, especially for difficult tasks. So the decoding time will be much longer.
Appendix B Some Problems of Other Long-context Benchmarks
We find existing long-context benchmarks problematic for our specific needs. For example:
  1. 1.
    LongBench Bai et al. (2023) is classic and widely used, but its average context length (most are under 16k) is relatively short for modern models, i.e., it is already too easy. And its data source is too old, leading to a risk of test data leakage. So we choose not to use it.
  2. 2.
    RulerHsieh et al. (2024) has controllable context length, but its tasks are almost all synthetic and most of them are needle-retrieval tasks, thus failing to reflect the model’s overall long-context capability in real-world scenarios.
  3. 3.
    LongBench-V2 Bai et al. (2024b) is well-designed to necessitate deep understanding and reasoning over very long context. But it is too challenging for 4B-level models (e.g., Qwen3-4B-Thinking only gets 35% accuracy, which is too close to the random guessing baseline of 25%), making the improvement of different methods less distinguishable. Moreover, since it is in a multiple-choice question format, the results may not be sufficiently reliable because the model has a 25% chance of guessing the correct option.
    However, despite these drawbacks, they have been widely used in long-context model benchmarking. Thus we still elect to conduct our experiments on LongBench V2 Bai et al. (2024b) and Ruler Hsieh et al. (2024) to verify the generalizability of our conclusions, as shown in Appendix C.
Appendix C Results of Other Models
We show the results of Qwen3-30B-A3B in Table 4, and the results of Llama3.1-8B-Instruct in Table 5. Due to the time-intensive nature of training, we only test a selective set of configurations with fine-tuning, thus the results are fewer. For Llama3.1-8B-Instruct on Ruler, the context length is controlled to 32k due to its weaker long-context capability. From the results, we observe that for any benchmark, the trend of accuracy change is mostly consistent with our previous conclusions, demonstrating their generalizability.
Appendix D Inference Efficiency
The TTFT, TPOT and total throughput when using vLLM are shown in Table 6. Since inference speed is highly dependent on hardware, implementation details, and workload characteristics, these numbers should be interpreted as reference values. However, from the results, we can still conclude that:
  1. 1.
    Interleaving Layers and FA Decode significantly slow down the speed compared to pure SWA.
  2. 2.
    Keep First $k$ Tokens has a negligible impact on efficiency.
  3. 3.
    Increasing the window size slightly increases inference time. For example, increasing from 2k to 4k decreases throughput by only 10%, but a 4k window generally achieves higher accuracy based on previous experiments. Therefore, in practice, a 4k window is a more common choice.
In theory, FA Decode should yield a decoding speed identical to that of full attention. Yet, in this table, we observe acceleration on TPOT. This is because vLLM-v1 typically mixes different requests’ prefilling and decoding tokens in one sequence to improve GPU utilization. Thus, the speeds of prefilling and decoding may affect each other. If processing only a single request, the situation differs. For example, when the generation length is set to 2000, we find decoding takes over 95% of the total time, rendering the acceleration of the prefilling stage negligible—i.e., SWA with FA Decode is almost unable to improve efficiency in such cases.
Appendix E Influence of Training Epochs
As shown in Table 7, training for more than 1 epoch yields no improvement. Therefore, we choose to train for only 1 epoch.
Appendix F Results of LightTransfer
LightTransfer Zhang et al. (2024) represents a promising attempt at SWA adaptation on full-attention models without pretraining. It proposes a layer selection method for SWA adaptation that calculates a "lazy ratio," represented by the ratio of attention from tokens at the end of the sequence (from a calibration dataset) to recent tokens versus global tokens. Layers with a higher "lazy ratio" are selected to apply SWA, while the rest retain full attention. This method is intuitive and theoretically sound, but our experiments reveal some negative results.
Since the complete code of LightTransfer is not open-source, we reproduce this method using LongAlign Bai et al. (2024a) as the calibration dataset for lazy layer detection, where the number of last tokens is set to 64, and the recent token window is set to 1024. From our experimental results shown in Table 8, we find that:
  1. 1.
    For Qwen3-4B, LightTransfer even yields a counterproductive effect; allowing lazy layers to use FA yields higher scores, while following the original method (letting non-lazy layers use FA) results in significantly lower scores.
  2. 2.
    For Qwen3-30B, it provides nearly no improvement over fixed-interval selection.
  3. 3.
    Only for Llama3.1-8B does LightTransfer show advantages.
Therefore, we conclude that LightTransfer does not yield stable performance across various models. Although fine-grained layer selection methods are theoretically superior, we believe they require further investigation before integration into our SWAA recipes.
Table 4: The accuracy of Qwen3-30B-A3B-Thinking and Qwen3-30B-A3B-Instruct on LongMemEval, LongBench-V2, and Ruler
No. SFT Window FA layers Keep FA LongMem LB-V2 Ruler
size first decode Think Non-T Think Non-T Think Non-T
0 False Full [] 0 False 79.2 71.6 49.7 42.6 95.6 99.4
1 False 2k [] 0 False 0.0 0.4 0.0 0.0 0.0 0.0
2 False 8k [] 0 False 0.0 0.2 0.0 0.0 0.0 0.0
3 False 2k [] 10 False 0.0 2.8 9.1 32.2 0.0 0.0
4 False 2k [] 0 True 0.2 0.2 0.0 0.0 0.0 0.0
5 False 2k [0, 2, 4, …] 0 False 21.0 28.4 20.1 25.8 1.4 17.0
6 False 2k [] 10 True 43.8 23.6 9.1 32.2 0.0 0.0
7 False 2k [] 100 True 58.6 22.2 10.4 28.2 2.4 3.0
8 False 2k [] 1k True 59.0 25.4 11.7 29.5 3.4 5.8
9 False 4k [] 10 True 49.8 26.6 26.8 30.9 3.0 1.2
10 False 2k [0, 2, 4, …] 10 True 74.8 63.0 22.1 33.6 85.4 98.8
11 False 2k [0, 4, 8, …] 10 True 48.8 23.8 12.4 29.5 1.6 1.8
12 False 2k [1, 3, 5, …] 10 True 51.6 24.0 30.2 28.9 17.8 64.2
13 False 2k [2, 6, 10, …] 10 True 64.8 44.2 21.1 35.6 35.8 50.2
14 False 4k [0, 2, 4, …] 100 True 74.6 64.4 29.5 35.9 89.4 99.0
15 True Full [] 0 False 79.6 72.0 43.6 42.0 97.6 99.4
16 True 2k [] 0 True 62.2 51.0 35.9 33.9 4.6 22.4
17 True 2k [] 100 True 65.6 50.8 36.6 32.9 10.8 57.8
18 True 2k [0, 2, 4, …] 0 True 72.6 $\backslash$ 41.3 $\backslash$ 91.0 $\backslash$
19 True 2k [0, 2, 4, …] 100 True 77.8 68.0 48.0 37.9 91.4 99.6
Table 5: The accuracy of Llama3.1-8B-Instruct on LongMemEval, LongBench-V2, and Ruler
No. SFT Window FA layers Keep FA LongMem LB-V2 Ruler
size first decode Non-Think Non-Think Non-Think
0 False Full [] 0 False 61.0 33.2 82.4
1 False 2k [] 0 False 0.6 0.0 0.0
2 False 8k [] 0 False 1.2 0.0 0.0
3 False 2k [] 10 False 1.8 28.9 0.0
4 False 2k [] 0 True 0.0 0.0 0.0
5 False 2k [0, 2, 4, …] 0 False 3.0 0.0 0.0
6 False 2k [] 10 True 16.8 28.9 1.0
7 False 2k [] 100 True 20.0 30.2 10.0
8 False 2k [] 1k True 24.2 30.2 18.4
9 False 4k [] 10 True 23.8 27.5 2.4
10 False 2k [0, 2, 4, …] 10 True 42.6 28.2 22.8
11 False 2k [0, 4, 8, …] 10 True 17.8 32.6 4.4
12 False 2k [1, 3, 5, …] 10 True 21.0 26.5 23.4
13 False 2k [2, 6, 10, …] 10 True 24.4 26.8 1.0
14 False 4k [0, 2, 4, …] 100 True 44.0 30.9 52.4
Table 6: Efficiency metrics of different SWAA configurations on vLLM. "FA layers = 1/4" means one fourth of total layers use full attention while the others use SWA.
window keep first FA decode FA layers TTFT (s) TPOT (s) Throughput (k tks/s)
Full 0 False None 1681.44 0.16 3.74
2k 0 False None 203.20 0.02 30.72
2k 100 False None 207.74 0.02 30.65
2k 0 False 1/2 938.00 0.09 6.70
2k 0 True None 963.39 0.11 6.39
2k 0 True 1/2 1321.39 0.14 4.72
2k 0 True 1/4 1141.66 0.12 5.43
4k 0 False None 233.07 0.02 27.03
4k 100 False None 237.87 0.02 26.74
4k 0 False 1/2 949.02 0.09 6.64
4k 0 True None 990.00 0.11 6.23
4k 0 True 1/2 1340.91 0.14 4.64
4k 0 True 1/4 1166.69 0.13 5.32
Table 7: Results of different training epochs of Qwen3-4B-Thinking on LongMemEval
SFT (epochs) window size FA layers keep first FA decode Acc
1 2k [] 0 True 58.0
2 2k [] 0 True 57.6
3 2k [] 0 True 56.0
Table 8: Results of LightTransfer on LongMemEval. "lazy" represents the half layers with higher lazy ratio, i.e. those which should apply SWA in theory. "non-lazy" represents the other part, i.e. those which should keep full attention.
SFT window size FA layers keep first FA decode Acc think Acc non-think
Model Group: Qwen3-4B
False 2k [0, 2, 4, …] 100 True 48.8 18.4
False 2k [1, 3, 5, …] 100 True 70.8 50.4
False 2k lazy 100 True 70.2 47.8
False 2k non-lazy 100 True 54.0 19.6
Model Group: Qwen3-30B-A3B
False 2k [0, 2, 4, …] 100 True 75.8 64.2
False 2k [1, 3, 5, …] 100 True 60.2 25.8
False 2k lazy 100 True 61.8 25.2
False 2k non-lazy 100 True 74.8 59.2
Model Group: Llama3.1-8B-Instruct
False 2k [0, 2, 4, …] 100 True $\backslash$ 39.8
False 2k [1, 3, 5, …] 100 True $\backslash$ 24.2
False 2k lazy 100 True $\backslash$ 20.2
False 2k non-lazy 100 True $\backslash$ 49.8